Hagar Industrial Systems Company (HISC) is trying to decide between two different conveyor belt systems. System A costs $290,000, has a 4-year life, and requires $93,000 in pretax annual operating costs. System B costs $370,000, has a 6-year life, and requires $87,000 in pretax annual operating costs. Both systems are to be depreciated straight-line to zero over their lives and will have zero salvage value. Whichever project is chosen, it will not be replaced when it wears out. The tax rate is 25 percent and the discount rate is 8 percent.

Calculate the NPV for both conveyor belt systems. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Ri9ngej7m9wdgCcDtiC1CA2AIUILYABYgtQAFiC1
Answer #1
System A:
Initial cost $ -2,90,000.00
PV of after tax annual operating costs = -93000*(1-25%)*(1.08^4-1)/(0.08*1.08^4) = $ -2,31,020.85
PV of tax shield on deprecitaion = (290000/4)*25%*(1.08^4-1)/(0.08*1.08^4) = $     60,032.30
NPV $ -4,60,988.55
System B:
Initial cost $ -3,70,000.00
PV of after tax annual operating costs = -87000*(1-25%)*(1.08^6-1)/(0.08*1.08^6) = $ -3,01,642.90
PV of tax shield on deprecitaion = (370000/6)*25%*(1.08^6-1)/(0.08*1.08^6) = $     71,269.39
NPV $ -6,00,373.50